OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

Slope intercept form is given by y = mx + b

OpenStudy (anonymous):

Given equation is 2x + 3y = 1200 convert this into y = mx + b form which is 3y = 1200 - 2x Where b (y - coordinate) is 1200 and slope m is -2x

OpenStudy (anonymous):

Could it be like this, though? : 3y= 2x +1200 , and then do (3/3)y = (-2/3) x + (1200/3)?

OpenStudy (anonymous):

Yes it can be same